OpenStudy (anonymous):

Solve the proportion. 2 5 -----= --- x-3 x 15 2 5 10

Nnesha (nnesha):

cross multiply :D \[\huge\rm \frac{ \color{Red}{a} }{ b } = \frac{\color{blue}{c}}{d}\] \[\large\rm \color{red}{a}d=\color{blue }{c}b\]

Nnesha (nnesha):

\[\huge\rm \frac{ \color{Red}{2} }{ x-3 } = \frac{\color{blue}{5}}{x}\] look at that example and multiply this

Nnesha (nnesha):

http://prntscr.com/8tb5ht first distribute the parentheses by 5 and then solve for x